A man is 6 feet tall and standing 80 feet from a flag pole with 35 degree incline. How tall is the flag pole?

Answer by stanbon(75887)

A man is 6 feet tall and standing 80 feet from a flag pole with 35 degree incline. How tall is the flag pole?
---
Draw the picture.
You have a right triangle on top of a rectangle.
Height of the triangle is "h".
Angle of inclination from the top of the peron's head to
the top of the flagpole is 35 degrees.
Base of the triangle is 80 ft.
--------
Equation:
tan(35) = h/80
h = 80*tan(35) = 56.02 ft
----
Total height of the flag pole is 6+56.02 = 62.02 ft.
==============================
Cheers,
Stan H.
================